OpenStudy (anonymous):

Solve by finding the value of x to the nearest degree. (it's an inverse cos problem and I'll post below in a comment...

OpenStudy (anonymous):

\[\cos^{-1} \sqrt{3}/3 = x\]

OpenStudy (anonymous):

calculator for this one

OpenStudy (anonymous):

if it was \[\cos^{-1}(\frac{\sqrt{3}}{2})\] you would get 60 immediately, but for this one you need a calculator

OpenStudy (anonymous):

ohh sorry i think it actually is over 2 my workbook had a mistake and i corrected it wrong...would you be able to do x=\[x=\tan^{-1} (-\sqrt{3}/3)\] ?